[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Mar 30, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Mar 30 16:12:05 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
Work continues on the NAWIPS RHEL upgrade. Cumulus at the surface desk was  
converted yesterday and is working well. The troublesome graphics cards in  
cumulus were upgraded as part of the conversion. NCO is working on the  
Atlantic desk today. Matt Sardi and Dave Zelinsky found and implemented a  
solution to web graphics formatting issue that was caused by the conversion  
to RHEL6.
HSU:
Once NCO is finished with the conversion of the TAFB N-AWIPS systems, they  
will do another pass through the HSU systems to address remaining issues.

AT&T is still troubleshooting the old Hurricane Hotline.
Other items:

NHC RFCs:
NCO is scheduled to make a change to our network on Tuesday, April 4th to  
fix the problem that is causing the slow download of the GFS download.

The GOES-16 data might be interrupted over the weekend due to maintenance  
in College Park. Monica is coordinating with the NWS on getting our data  
directly from our local dishes. This might be completed before the weekend,  
which would prevent the data loss.



-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170330/6645d889/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list